OpenStudy (anonymous):

If you had $10,000 and were investing it at 6% interest compounded continuously, how long would it take you to double your money, to the nearest tenth of a year?

OpenStudy (anonymous):

A = Pe^(rt) 20000 = 10000(e^.06x) x = 11.5 11.5 years.

OpenStudy (anonymous):

SWEET! Thank you very much..

OpenStudy (anonymous):

No Problem.

OpenStudy (anonymous):

did you get the steps?

OpenStudy (anonymous):

first of all doubling means doubling. it does not matter how much you start with, that is a red herring. just put \[2=e^{.06t}\] and solve for t

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

then take the log to get \[\ln(2)=.06t\] and so \[t=\frac{\ln(2)}{.06}\]

OpenStudy (anonymous):

ahhhh... yeah, ok.. cool!

OpenStudy (anonymous):

then use a calculator